OpenStudy (jameshorton):

Which is an example of an algebraic expression? A. g > 14 B. b + 32 C. m – 37 < 52 D. x + 4 = 13

OpenStudy (anonymous):

b

OpenStudy (jameshorton):

thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

hey im only in 6th grade so plz dont ask me for any 7th or up questions ok

OpenStudy (jameshorton):

im in 6th too

OpenStudy (jameshorton):

what school you go to

OpenStudy (itsbribro):

i'm in 9th if ya'll need help i can help you both maybe lol

OpenStudy (jameshorton):

Choose exactly two answers that are correct. 3. (3 pt) Which statement is true according to the order of operations? Choose exactly two answers that are correct. A. (11 + 3) • 2 = 14 • 2 = 28 B. 6 + 10 ÷ 2 = 6 + 5 = 11 C. 5 + (9 • 2) = 14 • 2 = 28 D. 8 – 4 ÷ 2 = 4 ÷ 2 = 2

OpenStudy (anonymous):

umm if you check my profile out it says

OpenStudy (itsbribro):

B and C

OpenStudy (anonymous):

ya B and C

OpenStudy (jameshorton):

Evaluate. (5 • 8 – 12)(36 ÷ 9 • 4)

OpenStudy (itsbribro):

try working it out PEMDAS

OpenStudy (anonymous):

or please excuse my dear aunt sally

OpenStudy (anonymous):

gtg

OpenStudy (itsbribro):

so what would be the first step?

OpenStudy (itsbribro):

@jameshorton

OpenStudy (jameshorton):

divide

OpenStudy (itsbribro):

Parenthesis Exponents Multiply Divide Add Subtract

OpenStudy (jameshorton):

parenthesis

OpenStudy (itsbribro):

so we would work out whats in the parenthesis first lets work from left to right so....

OpenStudy (itsbribro):

\[(5\times8-12)(36\div9\times4)\] whats 5 times 8

OpenStudy (itsbribro):

40 right?

OpenStudy (itsbribro):

\[(40-12)(36\div9\times4)\] so whats 40-12

OpenStudy (jameshorton):

28

OpenStudy (itsbribro):

good

OpenStudy (itsbribro):

now whats 36 divided by 9

OpenStudy (itsbribro):

how many times does 9 go into 36

OpenStudy (jameshorton):

0.25

OpenStudy (itsbribro):

no 9 goes into 36, 4 times

OpenStudy (anonymous):

ik the answer

OpenStudy (itsbribro):

so whats 4 times 4

OpenStudy (jameshorton):

16

OpenStudy (itsbribro):

okay now you have 28*16=

OpenStudy (jameshorton):

448

OpenStudy (itsbribro):

yes :) good job

OpenStudy (jameshorton):

Which options could be the next step in simplifying this expression? 15 + (5 + 21) Choose exactly two answers that are correct. A. 15 + 26 (use order of operations) B. 36 + 20 (use the commutative property) C. 20 + 36 (use the distributive property) D. 20 + 21 (use the associative property)

OpenStudy (itsbribro):

i think A and D

OpenStudy (itsbribro):

anything else?

OpenStudy (jameshorton):

yep

OpenStudy (jameshorton):

(3 pt) Which property does the statement illustrate? Column A Column B 1. (n + 6) • 7 = 7 • (n + 6) 2. 3 • (4 • 9) = (3 • 4) • 9 3. (2x • 5y) • 8 = 2x • (5y • 8) 4. 5 • 3 = 3 • 5 A. commutative property of multiplication B. associative property of multiplication

OpenStudy (jameshorton):

@dan815

OpenStudy (itsbribro):

i'm confused sorry lol

OpenStudy (jameshorton):

column b is post to be a and b

OpenStudy (jameshorton):

then you pick which one is accosicative or commutative

OpenStudy (mikeymaximum):

Are these multiple questions?

OpenStudy (itsbribro):

nvm i got it lol

OpenStudy (mikeymaximum):

Because if they are then that would make things 1,000,000,000,000,000,000,000,000,000,000,000,000,000,000,000,000,000,000x easier for me

OpenStudy (jameshorton):

do you know it

OpenStudy (mikeymaximum):

yeah if they are all multiple questions